I am stuck on correct answer for:
if the width of a rectangle is 3 inches less than its length and the perimeter of the rectangle is 94 inches, what is the length, in inches, of the rectangle?
I believe answer is 25 but I need to verify my work. Can you show me the breakdown to get correct resolve?

Answer by Nate(3500)

length = l
width = l - 3
Perimeter = 2(length + width)
94 = 2(l + l - 3)
94 = 2(2l - 3)
94 = 4l - 6
100 = 4l
25 = l
The length is 25 inches.